A single conducting lop of wire has an area of 7.2
Answer #1

From Faraday's law the induced emf in the coil is \varepsilon =-N\frac{\mathrm{d} \phi }{\mathrm{d} t}

\Rightarrow ixR=-Ax\frac{\left ( B_{2}-B_{1} \right )}{t}

0.32x110 = 7.2x10-2 x\frac{\mathrm{d} \phi }{\mathrm{d} t}

    \frac{\mathrm{d} \phi }{\mathrm{d} t}= 488.9T/s